# Positional Encoding in Transformer-Based Time Series Models: A Survey

### v1

@@ Discussion
Dataset-Specific Effectiveness
Dataset characteristics significantly influence encoding method performance, revealing important patterns for practical applications. In biomedical signals, particularly EEG datasets, TUPE and SPE consistently excel, suggesting these methods’ superior capability in capturing complex physiological patterns. Their effectiveness likely stems from their ability to model both local temporal dynamics and long-range dependencies characteristic of biological signals.
@@
Method Selection Guidelines
Our findings establish clear guidelines for method selection based on application characteristics, which we organize into three key considerations:
First, sequence length emerges as a primary factor in method selection. For sequences exceeding 100 time steps, the substantial performance gains of advanced methods (TUPE, SPE) clearly justify their increased computational complexity. However, for shorter sequences, the more modest performance improvements suggest that simpler methods may offer a better complexity-performance trade-off.
Second, data complexity significantly influences method effectiveness. Higher-dimensional data consistently benefits more from advanced encoding strategies, likely due to their superior ability to capture complex inter-dimensional temporal relationships. Similarly, binary classification tasks show larger relative improvements with sophisticated methods, possibly due to their enhanced ability to learn discriminative temporal features.
Third, computational requirements must be carefully considered. While advanced methods demonstrate superior performance, they also incur higher computational costs. This trade-off becomes more favorable for longer sequences where the performance gains are more substantial. For applications with real-time requirements or resource constraints, the selection of encoding methods should balance performance improvements against computational efficiency.
These insights provide valuable guidance for practitioners in selecting appropriate positional encoding methods based on their specific application requirements. The clear patterns in method effectiveness across different data characteristics enable informed decisions that optimize the performance-complexity trade-off for particular use cases.
Future Research Directions
Our analysis suggests several promising directions for future research. First, the development of adaptive encoding methods that automatically adjust their complexity based on sequence characteristics could provide optimal performance across varying data conditions. Second, investigating the integration of domain-specific temporal priors into encoding strategies could enhance performance for specific applications. Finally, exploring the relationship between positional encoding and attention mechanisms could lead to more efficient architectures for time series analysis.
Conclusion
@@
In this survey, we have presented a comprehensive analysis of positional encoding methods in transformer-based time series models, examining their effectiveness across various time series tasks and data types. Through extensive experimentation with eight distinct encoding methods across eleven diverse datasets, we demonstrated that advanced methods like TUPE and SPE consistently outperform traditional approaches while maintaining computational efficiency. Our findings establish clear relationships between sequence characteristics and encoding method effectiveness, providing valuable guidance for practitioners in selecting appropriate methods based on their specific application requirements. These insights lay the foundation for future research in developing more efficient and adaptive positional encoding strategies for time series analysis.

### v2

@@ 6.2 Architectural Considerations
6.3 Dataset-Specific Effectiveness
Dataset characteristics significantly influence encoding method performance, revealing important patterns for practical applications. In biomedical signals, particularly EEG datasets, TUPE and SPE consistently excel, suggesting these methods’ superior capability in capturing complex physiological patterns. Their effectiveness likely stems from their ability to model both local temporal dynamics and long-range dependencies characteristic of biological signals. The superior performance on biomedical signals reflects the methods’ ability to capture complex physiological feedback loops and temporal dependencies characteristic of biological systems [88, 89].
@@ 6.3 Dataset-Specific Effectiveness
6.4 Computational Considerations
The practical deployment of positional encoding methods requires careful consideration of computational trade-offs beyond accuracy metrics [90]. We analyze the computational characteristics of different PE methods to provide guidance for resource-constrained deployment scenarios.
Table 5: Training Time Analysis and Efficiency Metrics for PE methods on Melbourne Pedestrian Dataset in Transformer + Patch Embedding (100 epochs)
Method Time (s) Ratio Accuracy
Sin. PE 48.2 1.00 0.668
Learn. PE 60.1 1.25 0.702
RPE 128.4 2.66 0.724
tAPE 54.0 1.12 0.682
RoPE 67.8 1.41 0.690
eRPE 142.8 2.96 0.733
TUPE 118.3 2.45 0.745
ConvSPE 101.6 2.11 0.753
T-PE 134.7 2.79 0.742
AliBi 93.8 1.94 0.672
Figure 7: Performance vs computational complexity trade-off on Melbourne dataset. ConvSPE achieves the optimal efficiency frontier, delivering highest accuracy with moderate computational overhead.
[…]
6.5 Practical Guidelines for Method Selection
Based on our empirical analysis, we provide a systematic decision framework for positional encoding selection in transformer-based time series models.
6.5.1 Sequence Length-Based Recommendations
Short sequences ($L\leq 50$):
Use Learnable PE or tAPE. Advanced methods provide minimal gains (2–3%) that rarely justify computational overhead.
Medium sequences ($50<L\leq 100$):
SPE or eRPE recommended for 3–4% accuracy improvements. Consider T-PE when interpretability is required.
Long sequences ($L>100$):
TUPE for complex patterns, SPE for regular data, ConvSPE when linear complexity is essential. Advanced methods demonstrate 5–6% gains.
6.5.2 Domain-Specific Guidelines
  • Biomedical signals: TUPE >> SPE >> T-PE for physiological complexity
  • Environmental sensors: SPE >> eRPE for regular sampling patterns
  • High-dimensional data ($d>5$): Advanced methods consistently outperform simple approaches
6.5.3 Computational Resource Framework
Limited resources:
Sinusoidal PE, tAPE ($\mathcal{O}(Ld)$ complexity)
Balanced scenarios:
SPE, TUPE (optimal accuracy–efficiency trade-off)
Performance-critical:
TUPE, SPE regardless of computational cost
6.5.4 Architecture-Specific Considerations
For Time Series Transformers, prioritize content-position separation methods (TUPE) and relative positioning approaches (eRPE, SPE). For Patch Embedding Transformers, multi-scale approaches (T-PE, ConvSPE) handle hierarchical processing more effectively.
6.6 Limitations & Future Directions
While this survey provides a broad benchmark, certain limitations should be acknowledged, which in turn highlight avenues for future research. First, our study, though diverse with eleven datasets, cannot encompass the full spectrum of time series characteristics, such as extremely high-frequency data, highly non-stationary series with abrupt concept drifts, irregularly sampled data, or datasets with extensive missing values. The performance of PE methods on such specialized data types remains an area for further investigation. Second, our architectural analysis was focused on two representative transformer variants. The interplay and relative performance of PE strategies might differ with other emerging or highly specialized transformer backbones (e.g., those designed for ultra-long sequences or specific forecasting tasks). Third, while we benchmarked eight prominent PE methods, the field is dynamic, and newer or more niche techniques might offer different advantages. Finally, although we employed a consistent training protocol, an exhaustive hyperparameter search for every PE method across all datasets and architectures was beyond the scope of this survey; further fine-tuning could potentially lead to marginal shifts in relative performance for specific combinations.
These limitations inform future work: Developing adaptive PEs that adjust to data characteristics is a key direction. Further research into hybrid PE strategies and integrating domain-specific temporal priors could yield significant advances. Deeper interpretability analysis of PE effects on attention mechanisms is also needed. Finally, extending this benchmark to other tasks (e.g., forecasting, anomaly detection) and a more detailed computational overhead analysis would be valuable.
7 Conclusion
@@
In this survey, we have presented a comprehensive analysis of positional encoding methods in transformer-based time series models, examining their effectiveness across various time series tasks and data types. Through extensive experimentation with ten distinct encoding methods across eleven diverse datasets, we demonstrated that advanced methods like TUPE and SPE consistently outperform traditional approaches while maintaining computational efficiency. Our findings establish clear relationships between sequence characteristics and encoding method effectiveness, providing valuable guidance for practitioners in selecting appropriate methods based on their specific application requirements. These insights lay the foundation for future research in developing more efficient and adaptive positional encoding strategies for time series analysis.
Declarations
Conflicts of interest There are no conflicts of interest or competing interests.
Authors’ contributions
Habib Irani and Vangelis Metsis jointly conceptualized the study. Habib Irani conducted the literature review, implemented the experimental framework, and performed the benchmarking and analysis. Vangelis Metsis provided guidance on the methodology and contributed to the interpretation of results. Habib Irani wrote the initial manuscript draft, and Vangelis Metsis revised and edited the manuscript. All authors reviewed and approved the final version of the manuscript.